OpenStudy (anonymous):

The height of a cone is equal to three times the radius of the base. If the area 36pi cm^2, what is the volume of the cone?

OpenStudy (anonymous):

pi r^2=36pi, then, r=6 h=18 volume=1/3(36pi)x18=216picm^3